OpenStudy (anonymous):

A fantasy board game makes use of four dice. The first is a standard 6-faced die, the second die has 4 faces, and the other two have 10 faces each. If all the dice are thrown together and their combination dictates a certain outcome of the game, how many possible outcomes exist?

jimthompson5910 (jim_thompson5910):

1st die: 6 outcomes 2nd die: 4 outcomes 3rd die: 10 outcomes 4th die: 10 outcomes Each die is independent of one another, so we can multiply them out to get: 6*4*10*10 = 24*100 = 2400 There are 2,400 different outcomes for the 4 dice